# A countable-support symmetric iteration separating $\mathsf{PP}$ from $\mathsf{AC}$

### v5

@@ 5.6 What the iteration yields
Since $\mathcal{M}=\bigcup_{\lambda\in\mathrm{Ord}}M_{\lambda}$ (Theorem 5.59), there exists an ordinal $\lambda$ such that $w\in M_{\lambda}$. Fix a hereditarily symmetric $\mathbb{P}_{\lambda}$-name $\dot{w}\in\operatorname{HS}_{\lambda}$ with $\dot{w}^{G_{\lambda}}=w$.
@@
Choose set-sized hereditarily symmetric $\mathbb{P}_{\mathrm{Ord}}$-names $\dot{A},\dot{R}\in\operatorname{HS}_{\mathrm{Ord}}$ such that $A=\dot{A}^{G_{\infty}}$ and $R=\dot{R}^{G_{\infty}}$. By Lemma 5.67, letting $\alpha:=\sup(I(\dot{A})\cup I(\dot{R}))+1$, we have $A,R\in M_{\alpha}$.
@@
Let $x\in\mathcal{M}$. Choose a set-sized HS name $\dot{x}\in\operatorname{HS}_{\mathrm{Ord}}$ with $x=\dot{x}^{G_{\infty}}$. By Lemma 5.67, letting $\alpha:=\sup(I(\dot{x}))+1$, we have $x=\dot{x}^{G_{\alpha}}\in M_{\alpha}$. Since $M_{\alpha}\models\mathsf{SVC}(S)$ (by Lemma 5.74), there is an ordinal $\eta$ and a surjection $e:S\times\eta\twoheadrightarrow x$ in $M_{\alpha}$, hence also in $\mathcal{M}$. Thus $\mathcal{M}\models\mathsf{SVC}(S)$. The final statement follows from Fact 4.3. ∎
@@
Choose a set-sized hereditarily symmetric $\mathbb{P}_{\mathrm{Ord}}$-name $\dot{f}\in\operatorname{HS}_{\mathrm{Ord}}$ with $f=\dot{f}^{G_{\infty}}$. By Lemma 5.67, there is a set ordinal $\alpha$ such that $f\in M_{\alpha}$. If $f$ already has a right inverse in $M_{\alpha}$, we are done.
Otherwise, $f$ is eligible for scheduling by the $\mathsf{AC}_{\mathsf{WO}}$-bookkeeping clause in Definition 5.62. Hence by the guarantee clause in that definition, there exists a stage $\beta\geq\alpha$ such that either $f$ already has a right inverse in $M_{\beta}$, or else the successor iterand at stage $\beta$ includes the orbit $\mathsf{AC}_{\mathsf{WO}}$-package forcing $\mathbb{R}_{[f]}$. Assume the latter case and fix such $\beta$.
@@
Let $V\models\mathsf{ZFC}$. In the class forcing extension $V[G]$ obtained by the $\mathrm{Ord}$-length countable-support symmetric iteration over the Cohen-seed symmetric model $\mathcal{N}$ described above, the final symmetric model $\mathcal{M}$ satisfies
@@
$\mathsf{ZF}$ holds in $\mathcal{M}$ by Theorem 5.59. $\mathsf{DC}$ holds by Proposition 5.71. $\mathsf{PP}$ holds by Theorem 4.5 together with Propositions 5.68 and 5.77. $\mathsf{AC}_{\mathsf{WO}}$ holds by Proposition 5.77. Finally, $\neg\mathsf{AC}$ holds by Proposition 5.70. ∎
@@
By Theorem 4.5 together with Propositions 5.68 and 5.77, $\mathcal{M}\models\mathsf{PP}$. Hence there is an injection $i:X\hookrightarrow S\times\eta$.
@@
(2) $KWP_{1}$ using $\mathsf{PP}$. By Theorem 4.5 together with Propositions 5.68 and 5.77, $\mathcal{M}\models\mathsf{PP}$. Apply $\mathsf{PP}$ to the surjection $e:S\times\eta\twoheadrightarrow X$ to obtain an injection $i:X\hookrightarrow S\times\eta$. Composing with the injection $S\times\eta\hookrightarrow\mathcal{P}(\omega\times\eta)\cong\mathcal{P}(\omega\cdot\eta)$ from (1), we obtain $X\hookrightarrow\mathcal{P}(\theta)$ for $\theta=\omega\cdot\eta$. Thus $KWP_{1}$ holds in $\mathcal{M}$.

### v6

@@ 5.6 What the iteration yields
Since $\mathcal{M}=\bigcup_{\lambda\in\mathrm{Ord}}M_{\lambda}$ (Proposition 5.74), there exists an ordinal $\lambda$ such that $w\in M_{\lambda}$. Fix a hereditarily symmetric $\mathbb{P}_{\lambda}$-name $\dot{w}\in\operatorname{HS}_{\lambda}$ with $\dot{w}^{G_{\lambda}}=w$.
@@
Choose set-sized hereditarily symmetric $\mathbb{P}_{\mathrm{Ord}}$-names $\dot{A},\dot{R}\in\operatorname{HS}_{\mathrm{Ord}}$ such that $A=\dot{A}^{G_{\infty}}$ and $R=\dot{R}^{G_{\infty}}$.
By Lemma 5.73, there are ordinals $\alpha_{A},\alpha_{R}<\mathrm{Ord}$ such that $\dot{A}$, viewed as a $\mathbb{P}_{\alpha_{A}}$-name, belongs to $\operatorname{HS}_{\alpha_{A}}$ and $\dot{R}$, viewed as a $\mathbb{P}_{\alpha_{R}}$-name, belongs to $\operatorname{HS}_{\alpha_{R}}$, with
$A=\dot{A}^{G_{\alpha_{A}}}\qquad\text{and}\qquad R=\dot{R}^{G_{\alpha_{R}}}.$
Let $\alpha:=\max\{\alpha_{A},\alpha_{R}\}$. By Lemma 5.62, the canonical inclusions of $\dot{A}$ and $\dot{R}$ belong to $\operatorname{HS}_{\alpha}$, and their interpretations at stage $\alpha$ are still $A$ and $R$. Hence $A,R\in M_{\alpha}$.
@@
Let $x\in\mathcal{M}$. Choose a set-sized HS name $\dot{x}\in\operatorname{HS}_{\mathrm{Ord}}$ with $x=\dot{x}^{G_{\infty}}$. By Lemma 5.73, there is $\alpha<\mathrm{Ord}$ such that $\dot{x}$, viewed as a $\mathbb{P}_{\alpha}$-name, belongs to $\operatorname{HS}_{\alpha}$ and
$x=\dot{x}^{G_{\alpha}}\in M_{\alpha}.$
Since $M_{\alpha}\models\mathsf{SVC}(S)$ (by Lemma 5.81), there is an ordinal $\eta$ and a surjection $e:S\times\eta\twoheadrightarrow x$ in $M_{\alpha}$, hence also in $\mathcal{M}$. Thus $\mathcal{M}\models\mathsf{SVC}(S)$. The final statement follows from Fact 4.3. ∎
@@
Choose a set-sized hereditarily symmetric $\mathbb{P}_{\mathrm{Ord}}$-name $\dot{f}\in\operatorname{HS}_{\mathrm{Ord}}$ with $f=\dot{f}^{G_{\infty}}$. By Lemma 5.73, there is $\alpha<\mathrm{Ord}$ such that $\dot{f}$, viewed as a $\mathbb{P}_{\alpha}$-name, belongs to $\operatorname{HS}_{\alpha}$ and
$f=\dot{f}^{G_{\alpha}}\in M_{\alpha}.$
If $f$ already has a right inverse in $M_{\alpha}$, we are done.
By Lemma 5.68, there is a stage $\beta\geq\alpha$ such that the successor iterand at stage $\beta$ is the orbit $\mathsf{AC}_{\mathsf{WO}}$-package forcing $\mathbb{R}_{[f]}$. Fix such $\beta$.
@@
Let $V\models\mathsf{ZFC}$, and let $G\subseteq\mathbb{P}_{\mathrm{Ord}}$ be $V$-generic in the metatheoretic sense of Remark 2.1 for the $\mathrm{Ord}$-length countable-support symmetric iteration over the Cohen-seed symmetric model $\mathcal{N}$ described above. Then the final symmetric model $\mathcal{M}$ satisfies
@@
$\mathsf{ZF}$ holds in $\mathcal{M}$ by Theorem 5.64. $\mathsf{DC}$ holds by Proposition 5.78. $\mathsf{PP}$ holds by Theorem 4.5 together with Propositions 5.75 and 5.84. $\mathsf{AC}_{\mathsf{WO}}$ holds by Proposition 5.84. Finally, $\neg\mathsf{AC}$ holds by Proposition 5.77. ∎
@@
By Theorem 4.5 together with Propositions 5.75 and 5.84, $\mathcal{M}\models\mathsf{PP}$. Hence there is an injection $i:X\hookrightarrow S\times\eta$.
@@
(2) $KWP_{1}$ using $\mathsf{PP}$. By Theorem 4.5 together with Propositions 5.75 and 5.84, $\mathcal{M}\models\mathsf{PP}$. Apply $\mathsf{PP}$ to the surjection $e:S\times\eta\twoheadrightarrow X$ to obtain an injection $i:X\hookrightarrow S\times\eta$. Composing with the injection $S\times\eta\hookrightarrow\mathcal{P}(\omega\times\eta)\cong\mathcal{P}(\omega\cdot\eta)$ from (1), we obtain $X\hookrightarrow\mathcal{P}(\theta)$ for $\theta=\omega\cdot\eta$. Thus $KWP_{1}$ holds in $\mathcal{M}$.